OpenStudy (anonymous):

medal! Write the equation in slope-intercept from. What are the slope and y intercept? -6x+3y=12

OpenStudy (anonymous):

-6x+3y=12 ---isolate for y 3y=12+6x ----divide both sides by 3 y=(12+6x)/3 y=4+2x or y=2x+4 The slope-intercept form follows the format: y=mx+b, where m is your slope and b is your y-intercept. Therefore in y=2x+4, your slope is 2 and your y-intercept is 4.

OpenStudy (anonymous):

Slope defines how steep the line is running. The y-intercept is the value of y when x=0.

OpenStudy (anonymous):

Thanks so much! could you help me with another question?

OpenStudy (anonymous):

Sure, post it below and I'll give it a try.

OpenStudy (anonymous):

what is an equation of the line, in point slope from, that passes through the given point and has the giveslope? point (9,3) slope 4/9

OpenStudy (anonymous):

*given

OpenStudy (anonymous):

First, you need to calculate slope through the equation m=(y2-y1)/(x2-x1) Where y2 is the y-value of the second point, y1 is the y-value of the first point, and so on. Try that out first.

OpenStudy (anonymous):

dang it i broke my calculator

OpenStudy (anonymous):

so m=(0,3)/(0,9)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Oh, I'm sorry. I thought you had two points given to you. If you have the slope, then it's even easier. Use the format y=mx+b, substitute the slope (4/9) for m, and plug in the given point into the y and x of the equation. Solve for b.

OpenStudy (anonymous):

4/9=9+3?

OpenStudy (anonymous):

y=mx+b y=(4/9)x+b 9=(4/9)(3)+b 9-(4/9)(3)=b Solve for b. Once you have it, follow the format y=mx+b, and replace m with 4/9, and b with the number you found. And voila, you have the equation. (:

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

wait.... can you show me how to solve for b... sorry im so bad at math

OpenStudy (anonymous):

^Right up there. Rearrange the equation so that b is isolated on one side of the equal sign, and you get: 9-(4/9)(3)=b. Then just plug the left side into a calculator.

OpenStudy (anonymous):

ohhh thank you very much

OpenStudy (anonymous):

No problem (:

OpenStudy (anonymous):

So it gives me multiple choices though and it doesnt look like any of them

OpenStudy (anonymous):

Im guessing it would be y-9=4/9(x+4?

OpenStudy (anonymous):

What are the options?

OpenStudy (anonymous):

)*

OpenStudy (anonymous):

the options are: a. y-9=4/9(x+3) b. y-9=4/9(x-3) y+3=4/9(x-9) y-3=4/9(x-9)

OpenStudy (anonymous):

In point-slope form, the format is y=m(x-p)+q, where m is the slope and p,q is the point. The answer would be d, because p=9, and q=3 when you move it to the other side of the equal sign.

OpenStudy (anonymous):

Oh, now i understand.. Thank you i appreciate it

OpenStudy (anonymous):

I am going to post another question in a new thing if you dont mind looking at it.